OpenStudy (anonymous):

What are the steps to solve the following Differential Equation? y' = cos(2y-2x) Thank you in advance

OpenStudy (anonymous):

Any initial conditions?

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

Change it to\[\frac{ dy }{ dt }=\cos(2y)\cos(2x)-\sin(2y)\sin(2x)\]

OpenStudy (anonymous):

BTW, is this FOR differential equations class?

OpenStudy (anonymous):

I am studying for an exam and have come to this subject.Thank you for your response

OpenStudy (anonymous):

Is this for differential equations class though or for some different class?

OpenStudy (anonymous):

It's for a class in Applied Mathematics

OpenStudy (anonymous):

After DiffEQ, or before it?

OpenStudy (anonymous):

We are doing Differential Equations and Functions with Many Variables amongst other things in this class for the first time

OpenStudy (anonymous):

You're an undergrad stud right?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

What I would do is just use the equation above and then use two separate integrals and then since you're trying to undifferentiate X, just pull out the y terms and then integrate the X terms then you'll have your answer

OpenStudy (anonymous):

OK! Thank you very much!